# RSTeller: Scaling Up Visual Language Modeling in Remote Sensing with Rich Linguistic Semantics from Openly Available Data and Large Language Models

### v1

@@ Abstract
Abundant, well-annotated multimodal data in remote sensing are pivotal for aligning complex visual remote sensing (RS) scenes with human language, enabling the development of specialized vision language models across diverse RS interpretation tasks. However, annotating RS images with rich linguistic semantics at scale demands expertise in RS and substantial human labor, making it costly and often impractical. In this study, we propose a workflow that leverages large language models (LLMs) to generate multimodal datasets with semantically rich captions at scale from plain OpenStreetMap (OSM) data for images sourced from the Google Earth Engine (GEE) platform. This approach facilitates the generation of paired remote sensing data and can be readily scaled up using openly available data. Within this framework, we present RSTeller, a multimodal dataset comprising over 1 million RS images, each accompanied by multiple descriptive captions. Extensive experiments demonstrate that RSTeller enhances the performance of multiple existing vision language models for RS scene understanding through continual pre-training. Our methodology significantly reduces the manual effort and expertise needed for annotating remote sensing imagery while democratizing access to high-quality annotated data. This advancement fosters progress in visual language modeling and encourages broader participation in remote sensing research and applications. The RSTeller dataset is available at https://github.com/SlytherinGe/RSTeller.
@@ 1.1 Related works
The RS community has also made significant efforts to address the data scarcity problem for training VLMs for RS tasks. For readers’ easy comprehension of the scales of the related works, we compared them as in Fig. 1. Before the rise of large-scale pre-trained multimodal models like CLIP, several image-text datasets were developed, including UCM-Captions (Qu et al., 2016), Sydney-Captions (Qu et al., 2016), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), NWPU-Captions (Cheng et al., 2022), and RSVGD (Zhan et al., 2023). Among these, the largest dataset, NWPU-Captions, contains 31,500 images and 157,500 image-text pairs. The limited size of these datasets poses challenges for training VLMs like CLIP to perform well on RS tasks. To address this, (Liu et al., 2024a) compiled a RemoteCLIP dataset from three different types of RS tasks (segmentation, detection, and retrieval) across 17 individual datasets. They used hand-crafted rules to convert segmentation and detection annotations into multiple linguistic captions, resulting in a dataset with 165,745 images and 828,725 image-text pairs. Additionally, Zhang et al. (2023) proposed leveraging existing large-scale image-text datasets by filtering data related to RS and captioning several RS classification datasets with a pre-trained image caption model, creating an RS-related multimodal dataset called RS5M with 5 million image-text pairs. Training VLMs on these larger RS datasets has demonstrated improvements in their performance. However, there are limitations in building datasets from existing ones, as the upper bound is restricted by the quantity of available well-crafted datasets, which is limited compared to the vast Internet resources. In contrast, datasets like LAION (Schuhmann et al., 2021, 2022) were collected at the Internet scale, offering significantly larger amounts of data.
@@ 1.2 Contributions
In this work, we leverage openly available remote sensing data as an abundant resource for multimodal data generation. Specifically, we utilize the GEE platform to collect unlabeled aerial images, and the OSM platform as our semantic source. We employ Mixtral-7B (Jiang et al., 2024), an open-source LLM, to interpret plain OSM data. The OSM tag contents and their combinations are used to generate fluent, semantically rich descriptions for the aerial images. To ensure reproducibility and facilitate further scaling up of a dataset suitable for VLM training, we develop a comprehensive dataset generation workflow that encompasses raw data fetching, LLM-based captioning, and final dataset compilation.
Through this workflow, we introduce RSTeller, a multimodal dataset consisting of over one million remote-sensing images, each accompanied by multiple descriptive captions. We use the Measure of Textual Lexical Diversity (MTLD) score (McCarthy and Jarvis, 2010) as an indicator of the semantic richness of the captions. Our dataset is over two times more semantically rich than existing datasets (Fig. 1). Given the computational demands of LLM inference, our dataset provides a ready-to-use resource for researchers, eliminating the need for individual data generation.
Extensive experiments were conducted to validate the proposed methods and provide valuable insights for researchers training VLMs on RS data. First, we studied the relationship between the performance of the VLMs and the scale of RS domain data, highlighting the potential of further scaling up the training data. Then, we evaluated the effectiveness of RSTeller by continuing the pre-training of multiple VLMs and testing them on downstream tasks. The proposed data generation workflow and dataset are expected to be instrumental in advancing more sophisticated VLMs within the remote sensing field.
@@ 1.3 Paper structure
The structure of this paper is as follows: Section 2 details the proposed automatic data generation workflow. Section 3 describes and discusses the generated dataset, RSTeller, exploring its attributes in detail. Section 4 presents extensive experiments that explore the model performance concerning the data scale and the effectiveness of the proposed dataset. Section 5 provides some findings and suggestions according to our experiments. Section 6 outlines the limitations of the proposed work. Finally, Section 7 offers a summary of the research.
@@ 2.1.1 Raw Data Acquisition
The first process in the workflow is to fetch raw data from the GEE and OSM platforms into well-defined local databases. For ease of reference, we designate these databases as the ”raw data database,” as depicted in Process 1 of Fig. 2. This database is crucial for preserving the downloaded data and enabling asynchronous handling of different processes to accelerate the entire workflow. The GEE platform provides a vast collection of aerial images, each typically covering a wide range of geographical locations with tens or hundreds of thousands of pixels. The size of these images is too large for direct use in training a VLM, necessitating the cropping of images into smaller patches. GEE offers several flexible APIs for patch extraction and data download, from ee.Image.getThumbURL for JPEG or PNG formats to ee.Image.getDownloadURL for GeoTIFF or NumPy formats. For detailed API usage instructions, the official documentation [^3] https://developers.google.com/earth-engine/apidocs is recommended.
@@
The OSM data is another important source of semantic information for remote sensing. It can be queried via the Overpass API [^4] https://wiki.openstreetmap.org/wiki/Overpass_API using specific OverpassQL queries. OSM contains three types of geo-elements: nodes, ways, and relations. As nodes are points and too small to be seen in an RS image, we focus only on ways and relations. We implemented a two-step method to query OSM data to ensure the inclusion of all necessary geo-elements. The first step is to query all the ways and relations that intersect with or are within the image patches. This step includes most necessary geo-elements but may miss some large geometries that do not intersect but enclose the image patches, such as lakes or forests. The second step is to query all the ways and relations that enclose the center point of the patch, including those large geo-elements missed in the first step but are still very important to the image content.
The geographical contents of the image patches may evolve over time. Since the OSM data is not updated in real-time, we queried the OSM data for each image patch up to one month after it was captured by the sensor to best reflect the current image content. However, querying from the official endpoint [^5] https://overpass-api.de/api/interpreter requires following a strict usage policy, allowing fewer than 10,000 queries and 1 GB of data per day. Some third-party backends like Kumi Systems Overpass Instance [^6] https://overpass.kumi.systems/api/interpreter provide a free usage policy but may not guarantee the capacity and availability for big data processing due to server overload and network traffic. In our experience, using the Kumi endpoint can fetch OSM data for around 100,000 patches, each covering about 72,253.44 square meters in a day. Considering there could be patches not covered by OSM data and that captioning valid patches consumes much more time, it is inefficient to query the remote server each time a patch needs captioning. Therefore, it is better to cache the OSM data in a local database ahead of time or deploy the Overpass server locally. Using local OSM data during captioning significantly reduces the query time and improves the overall efficiency.
@@ 2.1.2 Raw Caption Generation
Once the data fetching process is underway and a sufficient amount of raw data has been acquired, the workflow proceeds to the generation of raw captions. This process involves using an LLM, specifically Mixtral-7B (Jiang et al., 2024) in this work, to interpret and describe the OSM data associated with each image patch according to predefined tasks. Instead of using an instruction fune-tuned LLM, like Mixtral-7B-instruct, the use of the base LLM is advantageous. It provides greater flexibility in generating captions because it is not constrained by specific instructional patterns. This allows it for better following the instructions and examples provided by the prompt, making the tone and style of the generated captions more consistent with the intended purpose.
In this process, a patch that has not yet been captioned is first sampled from the raw data database. The corresponding OSM data is then fetched from the cached OSM database. The OSM data undergoes processing based on our predefined tasks, which will be discussed in Subsection 2.2. This step checks whether the OSM data contains tags that can be used to generate captions for the patch. Patches that contain no relevant way or relation elements other than administrative borders, or where the OSM elements are too small to be visualized, are deemed unusable. These unusable patches are discarded and labeled as such in the raw data database. Only OSM data from valid patches is retained for further processing.
@@ 2.1.3 Caption Augmentation
Data augmentation in the linguistic aspect has been shown to enhance the training of vision language models (VLMs) (Fan et al., 2024). Unlike data augmentation in the visual domain, such as color jittering, noising, and random cropping, which can be easily implemented online during model training, augmenting the language aspect with consistent linguistic semantics and varied tones is resource-intensive, requiring the use of an LLM for high-quality results. Therefore, we incorporate a caption augmentation process into our workflow to generate multiple revisions of each caption.
During the caption augmentation process, captions that have not yet been revised are sampled from the annotation database. A specific revision task is defined to generate a prompt for the LLM based on the original caption and its initial task. The details of these tasks will be discussed in Subsection 2.2. The same LLM, Mixtral-7B, is used for this process. It takes the prompt as input and generates a revised caption for the image patch. The revised caption is then saved in the annotation database, and the status of the image patch is updated in the raw data database to reflect its completion.
This caption augmentation process is repeated until the desired number of revised captions is generated for each image patch. In RSTeller, our proposed dataset under this workflow, each image patch has at least two captions and up to five captions in total.
@@ 2.1.4 Dataset Compilation
During the inference of the LLM, some inevitable errors in the generated captions can lower the overall quality of the dataset. These errors can be categorized into two types: those that can be fixed and those that must be deleted. Fixable errors typically follow common patterns, such as repeated sentences within the same caption or undesired outputs appended after the caption ends, following some pattern of the input prompt. These redundant sentences and patterns can be removed through a heuristic process, allowing the remaining content to be used effectively. Unfixable errors usually include blank responses and duplicate captions for the same image patch. These errors cannot be corrected and must be deleted. The presence of these errors primarily stems from the imperfections of the LLM, and their frequency may be reduced through the evolution of the LLM, improvement of the prompts, or usage of a larger LLM.
@@ 2.2.1 Descriptive Tasks
The objective of the descriptive task is to generate detailed and contextually accurate descriptions of the OSM data associated with each image patch. These descriptions should capture the essential features and elements present within the patch, such as geographical landmarks, infrastructure, vegetation, and other notable features. These descriptions serve as the raw captions and are produced in Process 2. In a single image patch, there usually are a number of OSM elements, each with multiple attributes and tags. The information provided by all the OSM elements in an image patch can be overwhelming, making it challenging for an LLM to generate a concise and accurate description.
To streamline the task and enhance the efficiency of LLM annotation, we propose focusing on one distinctive OSM element per image patch for the description. In OSM, a two-dimensional object bounded by one or more linear ways and marked by appropriate tags is recognized as an ”area.” Following this approach, we categorize all OSM elements in an image patch into two groups: ”area” and ”non-area.” The ”area” category includes OSM elements like land use, natural features, and buildings, while the ”non-area” category includes elements not recognized as an ”area,” such as roads, coastlines, and other linear features. Given the significant differences in geometry and attributes between these two categories, we have designed two distinct tasks: Task 1 for the ”area” category and Task 2 for the ”non-area” category. Fig. 7 presents examples from the RSTeller dataset, showcasing image patches, selected OSM elements, and corresponding captions. The selected OSM elements are highlighted in red on the image patches. Image patches (2), (3), and (13) are captioned based on the ”non-area” category, while all other image patches are captioned based on the ”area” category.
@@
Distinctive area selection: In this step, the OSM element with the largest area within the image patch is selected as the most distinctive element to be described. The selected element undergoes a validation check to ensure it is not too small to be visible in the image. In this work, a threshold of 0.1 is used to filter out elements with an area smaller than this threshold portion of the image patch.
@@
Attribute interpretation: To enrich the semantic information of the selected element and facilitate the annotation process of the LLM, we extract several predefined attributes of the selected element and interpret them using a heuristic rule-based method. We have defined five attributes to describe the element, as shown in Table 1, including coarse location, shape, normalized size, simplified geometry, and is-cropped. The coarse location attribute identifies the center of the element within a $3\times 3$ grid, assigning a label to each grid cell (e.g., ”left-top”, ”center”, ”right-bottom”). The shape attribute describes the approximate shape of the element based on its geometry, defined by four categories: ”square”, ”rectangular”, ”circular”, and ”irregular”. These categories are determined by hard thresholding on the element’s area-to-perimeter ratio and shape-to-area ratio. The normalized size attribute indicates the relative size of the element compared to the image patch. The simplified geometry attribute represents the simplified and normalized geometry of the element, formatted as coordinate pairs enclosed in brackets (x, y). To avoid overwhelming the LLM annotators with extremely complex geometries, we use the Douglas-Peucker algorithm (Ramer, 1972) to simplify the element’s geometry. As some complex elements may be represented by multiple polygons, we encapsulate the coordinate pairs of each polygon within a list, and all polygons of an element are enclosed within braces. In this work, (0, 0) represents the bottom-left corner and (1, 1) the top-right corner, with all coordinates normalized to the range of [0, 1]. The is-cropped attribute is a boolean value that indicates whether parts of the element extend beyond the image patch, with values ”True” or ”False”. This attribute is determined by checking if the element visible in the image patch is cropped from a larger geometry.
@@
Coarse Location ”left-top”, ”center”, etc.
Shape ”square”, ”rectangular”, ”circular”, ”irregular”
Normalized Size ”0.300”, ”0.453”, etc.
Simplified Geometry ”{[(0.000, 0.500), (0.120, 0.400), …]}”, etc.
@@
Tag interpretation: OSM tags can sometimes be simple yet ambiguous. For example, the tag ”man_made: works” might be confusing to both the LLM and humans, as it can refer to various man-made structures. The OSM tag Wiki database provides explanations for such tags. According to the tag Wiki, ”man_made: works” refers to a factory or industrial production plant. This clarification removes ambiguity and makes the tag more meaningful for the LLM. Therefore, we query all OSM tag descriptions of the selected element from the tag Wiki. Using these descriptions, we interpret the tags into a uniform and linguistically consistent format, instead of merely using the key-value pairs. For tags with unbounded values, such as ”name: Jeff Memorial Highway,” we focus on interpreting the tag key instead. This results in a more comprehensive and accurate description of the selected element. An example of tag interpretation of the two kinds is shown in Table 2.
@@
man_made: works man_made: works. The tag belongs to the tag group ”NULL”. This tag means: ”A factory or industrial production plant”.
@@
name: Jeff Memorial Highway Its key is ”name”, which means ”the primary name: in general, the most prominent signposted name or the most common name in the local language(s).”. The tag belongs to a tag group ”names”. The tag value is Jeff Memorial Highway.

### v4

@@ Abstract
Abundant, well-annotated multimodal data in remote sensing are pivotal for aligning complex visual remote sensing (RS) scenes with human language, enabling the development of specialized vision language models across diverse RS interpretation tasks. However, annotating RS images with rich linguistic semantics at scale demands expertise in RS and substantial human labor, making it costly and often impractical. In this study, we propose a workflow that leverages large language models (LLMs) to generate multimodal datasets with semantically rich captions at scale from plain OpenStreetMap (OSM) data for images sourced from the Google Earth Engine (GEE) platform. This approach facilitates the generation of paired remote sensing data and can be readily scaled up using openly available data. Within this framework, we present RSTeller, a multimodal dataset comprising over 1.3 million RS images, each accompanied by two descriptive captions. Extensive experiments demonstrate that RSTeller enhances the performance of multiple existing vision language models for RS scene understanding through continual pre-training. Our methodology significantly reduces the manual effort and expertise needed for annotating remote sensing imagery while democratizing access to high-quality annotated data. This advancement fosters progress in visual language modeling and encourages broader participation in remote sensing research and applications. The RSTeller dataset is available at https://github.com/SlytherinGe/RSTeller.
@@ 1.1 Related Works
The RS community has also made significant efforts to address the data scarcity problem for training VLMs for RS tasks. To facilitate readers’ understanding of the semantic richness and scale of related works, we present a comparison in Fig. 1. This figure illustrates the semantic richness of each dataset proposed in previous works, as quantified by the Measure of Textual Lexical Diversity (MTLD) score (McCarthy and Jarvis, 2010), alongside the dataset scale, represented by both the number of textual tokens and the number of image-text pairs. Before the rise of large-scale pre-trained multimodal models like CLIP, several image-text datasets were developed, including UCM-Captions (Qu et al., 2016), Sydney-Captions (Qu et al., 2016), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), NWPU-Captions (Cheng et al., 2022), and RSVGD (Zhan et al., 2023). These datasets predominantly feature images paired with concise captions that provide basic descriptions of primary scenes or objects. While the RSICap dataset (Hu et al., 2023), introduced in the RSGPT work, sought to deliver richer semantic descriptions, its relatively low MTLD score highlights a limited diversity in lexical usage and sentence structures. Moreover, all the aforementioned datasets are constrained by scale, with NWPU-Captions being the largest, comprising 31,500 images and 157,500 image-text pairs. The limited size of these datasets poses challenges for training VLMs like CLIP to perform well on RS tasks.
To address the data scarcity problem, Liu et al. (2024a) compiled a RemoteCLIP dataset from three different types of RS tasks (segmentation, detection, and retrieval) across 17 individual datasets. They used handcrafted rules to convert segmentation and detection annotations into multiple linguistic captions, resulting in a dataset with 165,745 images and 828,725 image-text pairs. Additionally, Zhang et al. (2023) proposed leveraging existing large-scale image-text datasets by filtering data related to RS and captioning several RS classification datasets with a pre-trained image caption model, creating an RS-related multimodal dataset called RS5M with 5 million image-text pairs. Training VLMs on these larger RS datasets has demonstrated improvements in their performance. However, there are limitations in building datasets from existing ones, as the upper bound is restricted by the quantity of available well-crafted datasets, which is limited compared to the vast Internet resources. In contrast, datasets like LAION (Schuhmann et al., 2021, 2022) were collected at an Internet scale, offering significantly larger amounts of data.
@@ 1.2 Contributions
In this work, we leverage openly available remote sensing data as an abundant resource for multimodal data generation. Specifically, we utilize the GEE platform to collect unlabeled aerial images, and the OSM platform as our semantic source. We employ two variant open-source LLMs to interpret plain OSM data. The OSM tag contents and their combinations are used to generate fluent, semantically rich descriptions for the aerial images. To ensure reproducibility and facilitate further scaling up of a dataset suitable for VLM training, we develop a comprehensive dataset generation workflow that encompasses raw data fetching, LLM-based captioning, and final dataset compilation.
Through this workflow, we introduce RSTeller, a multimodal dataset consisting of over 1.3 million remote-sensing images, each accompanied by two descriptive captions. The semantic richness of these captions, measured using the MTLD score (McCarthy and Jarvis, 2010), is about twice that of existing datasets (Fig. 1). Given the computational demands of LLM inference, our dataset provides a ready-to-use resource for researchers, eliminating the need for individual data generation.
Extensive experiments were conducted to validate the proposed methods and provide valuable insights for researchers training VLMs on RS data. First, we demonstrated the effectiveness of RSTeller by the continual pre-training of multiple VLMs and evaluating their performance on downstream tasks. Second, we performed an ablation study to analyze the impact of data properties on model training. Third, we examined the relationship between VLM performance and the scale of RS-specific training data, underscoring the benefits of scaling data further. The proposed data generation workflow and dataset are expected to be instrumental in advancing more sophisticated VLMs within the remote sensing field.
@@ 1.3 Paper Structure
The structure of this paper is as follows: Section 2 details the proposed automatic data generation workflow. Section 3 describes and discusses the generated dataset, RSTeller, exploring its attributes in detail. Section 4 presents experiments evaluating the effectiveness of the proposed dataset, an ablation study on data properties, and an analysis of VLM performance with respect to data scale. Section 5 provides some findings and suggestions according to our experiments. Section 6 outlines the limitations of the proposed work. Finally, Section 7 offers a summary of the research.
@@ 2.1.1 Raw Data Acquisition
The first process in the workflow is to fetch raw data from the GEE and OSM platforms into well-defined local databases. For ease of reference, we designate these databases as the “raw data database,” as depicted in Process 1 of Fig. 2. This database is crucial for preserving the downloaded data and enabling asynchronous handling of different processes to accelerate the entire workflow. The GEE platform provides a vast collection of aerial images, each typically covering a wide range of geographical locations with tens or hundreds of thousands of pixels. The sizes of these images are too large for direct use in training a VLM, necessitating the cropping of images into smaller patches. GEE offers several flexible APIs for patch extraction and data download, from ee.Image.getThumbURL for JPEG or PNG formats to ee.Image.getDownloadURL for GeoTIFF or NumPy formats. For detailed API usage instructions, the official documentation [^3] https://developers.google.com/earth-engine/apidocs is recommended.
@@
The OSM data is another important source of semantic information for remote sensing. It can be queried via the Overpass API [^4] https://wiki.openstreetmap.org/wiki/Overpass_API using specific OverpassQL queries. OSM contains three types of geo-elements: nodes, ways, and relations. As nodes are points and too small to be seen in an RS image, we focus only on ways and relations. A notable challenge in utilizing the Overpass API is that it processes ways and relations as collections of nodes. When using the bounding box filter, one of the most common approaches to querying spatial data provided by Overpass, only geometries with at least one node located within the bounding box are retrieved. As a result, large enclosing geometries, such as lakes or forests, that fully contain the bounding box but lack any nodes within it are not captured. To address this limitation, we implemented a two-step method to ensure comprehensive inclusion of relevant geo-elements. In the first step, we queried ways and relations using the bounding box filter to retrieve all geometries that satisfy the filter’s policy. This step captures many essential geo-elements but may miss larger enclosing geometries, as described. In the second step, we performed an additional query using the is_in operator, specifically targeting geometries tagged as “area,” to identify elements that enclose the center point of the image patch. This approach ensures the inclusion of large geo-elements that might have been overlooked in the first step, thereby providing a more complete representation of the image content.
The geographical contents of the image patches may evolve over time. Since the OSM data is not updated in real-time, we queried the OSM data for each image patch up to one month after it was captured by the sensor to best reflect the current image content. However, querying from the official endpoint [^5] https://overpass-api.de/api/interpreter requires following a strict usage policy, allowing fewer than 10,000 queries and 1 GB of data per day. Alternative third-party backends, such as the Kumi Systems Overpass Instance [^6] https://overpass.kumi.systems/api/interpreter, offer more lenient usage policies but may be affected by server overload and network traffic. Based on our experience, the Kumi endpoint enabled the retrieval of OSM data for approximately 100,000 patches daily, each covering an area of 72,253.44 square meters. Given the computational demands of captioning valid patches and the potential for incomplete OSM coverage, querying remote servers on-demand was deemed inefficient. Caching OSM data locally prior to captioning or deploying a local Overpass server can significantly reduce query times and enhance overall efficiency. In this study, we opted to cache data from the Kumi endpoint due to the technical challenges and resource requirements associated with setting up a local Overpass server. By caching sufficient data in advance, we ensured uninterrupted captioning without the need for additional infrastructure. For larger-scale projects with sufficient computational resources and expertise, deploying a local Overpass server remains a recommended approach to further optimize scalability and performance.
@@ 2.1.2 Raw Caption Generation
Once the data fetching process is underway and a sufficient amount of raw data has been acquired, the workflow proceeds to the generation of raw captions. This process uses LLMs, specifically Mixtral-Nemo-Instruct (12B) and Mixtral-Small-Instruct (22B) (Jiang et al., 2024), to interpret and describe the OSM data for each image patch based on predefined tasks randomly. Model size was chosen according to the balance between caption quality and computational cost. The primary focus of our task is rephrasing raw materials into fluent sentences, with some minor reasoning to infer the surroundings and appearance of OSM elements. Since this reasoning is relatively simple compared to tasks like solving complex mathematical problems or multi-step logical reasoning, smaller models are sufficient for our needs. They provide high-quality outputs while minimizing computational costs, making them ideal for large-scale dataset generation. The Mixtral family was selected for its strong performance across a range of natural language tasks, accessibility, and flexible use policy. Additionally, using models of different sizes and versions introduces diversity in the captions, which could benefit future research on model variation and its impact on generated data.
In this process, a patch that has not yet been captioned is first sampled from the raw data database. The corresponding OSM data is then fetched from the cached OSM database. The OSM data undergoes processing based on our predefined tasks, which will be discussed in Subsection 2.2. This step checks whether the OSM data contains tags that can be used to generate captions for the patch. Patches that contain no relevant way or relation elements other than administrative borders, underground features, or where the OSM elements are too small to be visualized, are deemed unusable. These unusable patches are discarded and labeled as such in the raw data database. Only OSM data from valid patches are retained for further processing.
@@ 2.1.3 Caption Augmentation
Data augmentation in the linguistic aspect has been shown to enhance the training of VLMs (Fan et al., 2024). Unlike data augmentation in the visual domain, such as color jittering, noising, and random cropping, which can be easily implemented online during model training, augmenting the language aspect with consistent linguistic semantics and varied tones is resource-intensive, requiring the use of an LLM for high-quality results. Therefore, we incorporate a caption augmentation process into our workflow to generate one revision of each caption.
During the caption augmentation process, captions that have not yet been revised are sampled from the annotation database. A specific revision task is defined to generate a prompt for the LLM based on the original caption and its initial task. The details of these tasks will be discussed in Subsection 2.2. The same LLMs, Mixtral-Nemo-Instruct and Mixtral-Small-Instruct, are used for this process. It takes the prompt as input and generates a revised caption for the image patch. The revised caption is then saved in the annotation database, and the status of the image patch is updated in the raw data database to reflect its completion.
[…]
@@ 2.1.4 Dataset Compilation
During the inference of the LLM, certain errors in the generated captions are inevitable and can compromise the overall quality of the dataset. These errors are broadly categorized into two types: fixable errors and unfixable errors. Fixable errors typically follow recognizable patterns, such as repeated sentences appended after the caption or undesired outputs generated due to specific patterns in the input prompt. To address these, we implemented a systematic process. First, captions are segmented into individual sentences, and the uniqueness of each sentence is verified to detect and remove duplications. Second, common patterns of undesired outputs are identified through a comprehensive survey of the dataset. Based on these observations, we developed and applied a set of regular expressions to automatically detect and eliminate such errors, ensuring that the retained captions are concise and relevant. Unfixable errors, such as blank responses, invalid symbols, unexpected outputs, or duplicate captions generated for the same image patch, cannot be corrected and must be removed from the dataset. Similar to the handling of fixable errors, we analyzed the dataset to identify overarching patterns of unfixable errors and utilized handcrafted regular expressions to detect and eliminate them efficiently. The occurrence of these errors is primarily attributed to the limitations of the LLM, such as sensitivity to prompt design or variability in generative consistency. Moreover, error patterns may differ across various LLMs due to differences in their training data and architectures. However, the frequency of errors can be significantly reduced through several strategies: refining input prompts to provide clearer guidance or employing more advanced LLMs with enhanced capabilities. As LLM technology continues to advance, these measures are anticipated to further enhance the reliability and semantic accuracy of generated captions, thereby consistently elevating the quality of synthetic data produced by LLMs.
@@
Fig. 3: Inspection of the RSTeller dataset, showcasing four image patches with corresponding OSM tags, captions, and their revisions. The map highlights sample locations, while the figure illustrates OSM area and non-area elements with masked boundaries or lines. Crossed-out tags indicate filtered elements excluded from captioning. Captions represent results from Task 1 or Task 2, while revisions are outcomes of Task 3.
@@ 2.2.1 Descriptive Tasks
The objective of the descriptive task is to generate detailed and contextually accurate descriptions of the OSM data associated with each image patch. These descriptions should capture the essential features and elements present within the patch, such as geographical landmarks, infrastructure, vegetation, and other notable features. These descriptions serve as the raw captions and are produced in Process 2.
In a single image patch, there usually are a number of OSM elements, each with multiple attributes and tags. The information provided by all the OSM elements in an image patch can be overwhelming, making it challenging for an LLM to generate a concise and accurate description. To streamline the task and reduce the potential for hallucination, we define a focused task that captions only one distinctive OSM element per image patch. By reducing the input context length and concentrating on a single element, we ensure that the LLM processes relevant information without being overwhelmed by extraneous details. This approach simplifies the generation process and enhances the semantic accuracy of the captions.
In OSM, a two-dimensional object bounded by one or more linear ways and marked by appropriate tags is recognized as an “area”. The simplest areas are defined by closed ways, having the same start and end node. More complex areas are defined using multipolygon relations, which may specify multiple boundary ways. Following this approach, we categorize all OSM elements in an image patch into two groups: “area” and “non-area”. The “area” category includes OSM elements like land use, natural features, and buildings, while the “non-area” category includes elements not recognized as an “area,” such as roads, coastlines, and other linear features. Given the significant differences in geometry and attributes between these two categories, we have designed two distinct tasks: Task 1 for the “area” category and Task 2 for the “non-area” category. Fig. 3 presents examples from the RSTeller dataset, showcasing image patches, selected OSM elements, and corresponding captions. The selected OSM elements are highlighted in red on the image patches. Image patches (a) and (b) are captioned based on the “area” category, while the others are captioned based on the “non-area” category.
@@
Distinctive area selection: In this step, OSM elements smaller than 0.05 of the image patch area are filtered out to exclude insignificant features. From the remaining elements, if more than three candidates exist, one is randomly selected from the top three by area; otherwise, selection is made from all available elements. This approach highlights spatially dominant features while enhancing diversity and ensuring broader geo-element representation across image patches.
@@
Attribute interpretation: To enable robust and context-aware training for VLMs, we carefully selected attributes that encapsulate the critical geometric and semantic properties of geo-elements. These attributes ensure the dataset effectively captures essential spatial, morphological, and contextual details, enhancing the descriptive and inferential capabilities of VLMs. By striking a balance between granularity and interpretability, they optimize the dataset’s utility across diverse remote sensing tasks. As summarized in Table 1, we define five key attributes: coarse location, shape, normalized size, simplified geometry, and is-cropped. The coarse location attribute identifies the center of the element within a $3\times 3$ grid, assigning a label to each grid cell (e.g., “left-top,” “center,” “right-bottom”). This attribute provides spatial context, helping VLMs map object positions relative to the image patch and supporting tasks requiring spatial reasoning. The shape attribute describes the approximate shape of the element based on its geometry, defined by four categories: “square,” “rectangular,” “circular,” and “irregular”. This simplifies the representation of complex geometries, enabling models to recognize and differentiate structural patterns efficiently. These categories are determined by hard thresholding on the element’s area-to-perimeter ratio and shape-to-area ratio. The normalized size attribute quantifies the relative scale of the element within the image patch, offering essential context for understanding the prominence and spatial importance of the element in the scene. The simplified geometry attribute encodes the spatial outline of the element in a compact, normalized format, preserving essential geometric characteristics while minimizing complexity. Although not directly included in the final captions, this attribute functions as supplementary input for the captioning LLM, enhancing its interpretation of the size, shape, and location attributes to produce more informative and contextually accurate descriptions. To avoid overwhelming the LLM annotators with extremely complex geometries, we use the Douglas-Peucker algorithm (Ramer, 1972) to simplify the element’s geometry. As some complex elements may be represented by multiple polygons, we encapsulate the coordinate pairs of each polygon within a list, and all polygons of an element are enclosed within braces. In this work, (0, 0) represents the bottom-left corner and (1, 1) the top-right corner, with all coordinates normalized to the range of [0, 1]. The is-cropped attribute indicates whether parts of the element extend beyond the image patch. This provides critical information about spatial boundaries, enhancing the model’s ability to handle cropped or incomplete objects—a common scenario in remote sensing datasets. This attribute is determined by checking if the element visible in the image patch is cropped from a larger geometry.
In some cases, an area element can be represented as a multipolygon. This typically occurs due to complex geometries or when a larger geometry is cropped into separate parts. Fig. 3(b) illustrates an example where a farmland area is split into two separate polygons. In such cases, we provide a coarse location interpretation for each polygon but focus on describing the shape of the largest one. The normalized size is calculated based on the total area of the entire multipolygon. Attributes such as simplified geometry and is-cropped are also derived from the overall multipolygon structure.
@@
Coarse Location “left-top”, “center”, etc.
Shape “square”, “rectangular”, “circular”, “irregular”
Normalized Size “0.300”, “0.453”, etc.
Simplified Geometry “{[(0.000, 0.500), (0.120, 0.400), …]}”, etc.
@@
Tag interpretation: OSM tags serve as the primary semantic source for the captioning task but can also introduce significant noise, which may adversely affect the performance of captioning models and lead to hallucinations. This noise primarily arises from two sources: irrelevant tags and ambiguous tags.
Irrelevant tags, such as “TIGER,” “website,” and other metadata fields, often contain ID digits or information unrelated to the visual features of the image. To mitigate this issue, we apply a rigorous tag filtering process using predefined regular expressions (see A for details). This filtering ensures that only meaningful and relevant data is passed to the LLM for caption generation, thereby minimizing the risk of hallucinations. The effectiveness of this filtering is illustrated in Fig. 3, where unused tags are visually marked with strikethroughs.
[…]
Ambiguity in OSM tags further complicates their direct use in captioning, as some tags can be vague and difficult for both LLMs and human annotators to interpret. For instance, the tag “man_made: works” might refer to various man-made structures, but according to the OSM Wiki, it specifically denotes a factory or industrial production plant. To resolve such ambiguities and enhance the semantic clarity of the tags, we query the OSM Wiki for all tag descriptions corresponding to the selected elements. These descriptions are reformatted into a uniform and linguistically consistent structure, replacing the raw key-value pairs as input to the LLM. For tags with unbounded values, such as “name: Jeff Memorial Highway,” we focus on interpreting the tag key (e.g., “name”) rather than the specific value. This standardized and enriched representation ensures a comprehensive and accurate description of each selected element. An example of these two types of tag interpretation is provided in Table 2.
[…]
Ambiguity in OSM tags further complicates their direct use in captioning. For instance, the tag “man_made: works” can be vague and difficult for both LLMs and human annotators to interpret, as it might refer to a variety of man-made structures. To address this, we leverage the OSM tag Wiki database, which provides detailed descriptions of tag meanings. According to the Wiki, “man_made: works” specifically refers to a factory or industrial production plant. By incorporating these explanations, we resolve ambiguities and enhance the semantic clarity of the tags.
[…]
To standardize and enrich the input for the LLM, we query the OSM Wiki for all tag descriptions corresponding to the selected elements. These descriptions are reformatted into a uniform and linguistically consistent structure, replacing the raw key-value pairs. For tags with unbounded values, such as “name: Jeff Memorial Highway,” we focus on interpreting the tag key (e.g., “name”) instead of the value. This approach ensures a comprehensive and accurate description of each selected element. An example of these two types of tag interpretation is provided in Table 2.
@@
man_made: works man_made: works. The tag belongs to the tag group “NULL”. This tag means: “A factory or industrial production plant”.
@@
name: Jeff Memorial Highway Its key is “name”, which means “the primary name: in general, the most prominent signposted name or the most common name in the local language(s).”. The tag belongs to a tag group “names”. The tag value is Jeff Memorial Highway.